OpenStudy (anonymous):

da-chun has five times as many coins as austin. Da-chun also has 16 more than austin

OpenStudy (anonymous):

let da-chun's amount of coins =d. let austin's amount of coins=a. 5a=d. d-16=a. 5a-16=a. 4a=16, therefore a=4. Austin has 4 coins, da-chun has 20.

OpenStudy (anonymous):

A slightly different approach.\[\left\{\frac{d}{a}\text{ = }5,d-16=a\right\}\]